OpenStudy (anonymous):

Emergency help on an Algebra II exam! Prove: sin θ - sin θ•cos2 θ = sin3 θ. You must show all work.

OpenStudy (nincompoop):

it looks like this is an exam

OpenStudy (anonymous):

It's a practice exam, but it's worth a good bit of points and while I can understand the proofs written out in my lessons, this particular problem is stumping me. I've been working on these tests for hours :(

OpenStudy (anonymous):

So I know that identity of sin^2theta + cos^2theta = 1, or...something pretty close to that...

OpenStudy (nincompoop):

show me the attempts you have made.

OpenStudy (anonymous):

Okay. So the sin theta should cancel, right? cos^2 theta = sin^3 theta ****(that 3 is a cube, not sin(3))**** I don't know what to do with that!

OpenStudy (nincompoop):

how does it cancel?

OpenStudy (anonymous):

Bc it's just subtracted. Like 1 - 1 = 0.

OpenStudy (anonymous):

sin theta - sin theta = 0.

OpenStudy (nincompoop):

no

OpenStudy (nincompoop):

sin is multiplied to cos sin θ - (sin θ•cos2 θ) = sin3 θ ORDERS of operation

OpenStudy (anonymous):

Parentheses first, then

OpenStudy (nincompoop):

btw, is that \(cos^2 \theta \) or \(cos (2 \theta) \)

OpenStudy (anonymous):

The first, it's squared.

OpenStudy (anonymous):

Okay, so...then do we distribute the negative into the parentheses??

OpenStudy (anonymous):

I'm thinking maybe we subtract sin theta from both sides to give us (sin theta * cos^2 theta) = sin^2 theta

OpenStudy (nincompoop):

is that \(sin^3 \theta \) or \(sin (3 \theta) \)

OpenStudy (anonymous):

Cubed, the first one.

OpenStudy (nincompoop):

you need to start typing these things appropriately because they mean different things

OpenStudy (anonymous):

I'm sorry, I didn't realize it before I posted bc I just copy and pasted it to preserve the theta symbols :(

OpenStudy (nincompoop):

like this? PROVE \(sin ~\theta - (sin~ \theta ~cos^2 \theta) = sin^3 \theta \)

OpenStudy (anonymous):

There weren't any parentheses, but otherwise, yes.

OpenStudy (nincompoop):

so tell me what should be the best identity we can use?

OpenStudy (anonymous):

The only one I know of is the one I typed above, the sin^2 theta + cos^2 theta = 1.

OpenStudy (anonymous):

We didn't focus a lot on identities, so I suppose I may be forgetting some, but they aren't in my notes =/

OpenStudy (nincompoop):

what did you focus on?

OpenStudy (nincompoop):

let us use trigonometric expansion and make use of some algebra techniques like factoring, yes?

OpenStudy (nincompoop):

do you think you can factor the left-hand-side (LHS) of the equation for me?

OpenStudy (anonymous):

Sorry! Lost connection!

OpenStudy (anonymous):

I can tryyy, hold on

OpenStudy (nincompoop):